OpenStudy (anonymous):

What is the area of the triangle? A. 504 cm² B. 648 cm² C. 1008 cm² D. 1296 cm²

OpenStudy (flvs.net):

I want to say A. :/

OpenStudy (hybrik):

@Flvs.net explain more, the formula of a triangle is A = 1/2 BH

OpenStudy (flvs.net):

A = 504

OpenStudy (hybrik):

Do 72*18 then divide that by 2

OpenStudy (flvs.net):

Me?

OpenStudy (anonymous):

yea

OpenStudy (hybrik):

Which is 648

OpenStudy (hybrik):

A = 648 units squared

OpenStudy (flvs.net):

That's not the answer

OpenStudy (hybrik):

Explain

OpenStudy (flvs.net):

You are wrong

OpenStudy (anonymous):

explain

OpenStudy (hybrik):

Explain how

OpenStudy (hybrik):

I used the base and height, you used the hypothenuse, which is clearly wrong

OpenStudy (flvs.net):

Because if you have this when you start out (one sec.)

OpenStudy (anonymous):

k

OpenStudy (flvs.net):

\[A = \frac{ h _{b}b }{ 2 }\]

OpenStudy (flvs.net):

if you start out with that

OpenStudy (flvs.net):

nevermind he's gone

OpenStudy (anonymous):

@hybrik

OpenStudy (flvs.net):

I was gonna do it but forget it

OpenStudy (anonymous):

@hybrik

OpenStudy (flvs.net):

It's fine.

OpenStudy (hybrik):

Yes?

OpenStudy (hybrik):

the real formula is 1/2 BH to find the area of the triangle

OpenStudy (flvs.net):

For the base I put in 56

OpenStudy (hybrik):

That is incorrect, that is the hypothenuse

OpenStudy (flvs.net):

and the Height \[(h _{b})\]

OpenStudy (flvs.net):

18

OpenStudy (anonymous):

i know somone @Mehek14

Mehek (mehek14):

wut?

OpenStudy (flvs.net):

and got \[504 cm ^{2}\]

OpenStudy (anonymous):

what is the answer ?

OpenStudy (flvs.net):

Whatever i'm done

OpenStudy (anonymous):

ok

Mehek (mehek14):

648

OpenStudy (flvs.net):

How?

Mehek (mehek14):

@hybrik was correct

OpenStudy (hybrik):

@Flvs.net Being honest your answer is incorrecct.

OpenStudy (flvs.net):

Whatever, i'm freakin stupid

OpenStudy (flvs.net):

Sorry @speartonion

Mehek (mehek14):

\(\dfrac{1}{2}*72*18\) @Flvs.net no you're not everyone makes mistakes just learn from them 72 * 18 = 1926 \(\dfrac{1926}{2}=648\)

OpenStudy (anonymous):

YOUR FREAKING AWESOME @Mehek14

OpenStudy (anonymous):

and she is right @Flvs.net your not

OpenStudy (flvs.net):

I missed two in a row.

OpenStudy (flvs.net):

and you trusted i could it.